OpenStudy (alexandrad):

Write the equation of the line that passes through (3, 4) and (2, −1) in slope-intercept form.

OpenStudy (alexandrad):

A. y=3x-7 B. y=3x-5 C. y=5x-11 D. y=5x-9

OpenStudy (zyi6):

So, first we have to find the slope of the line: Remember that the formula to find slope is \[y _{1} - y _{2} / x _{1} - x _{2}\] So, we can substitute the given values in to find the slope: 4 -(-1) / 3 - 2 = 5 So now that we have our slope, we can start our equation like so: y = 5x + b So, now we need our y-intercept. In order to do this, we can substitute in one of our points into the equation to solve for the y-intercept or the b of the equation. It does not matter which point you choose: I will use (3, 4) here: 4 = 5(3) + b Then solve for b: 4 = 15 + b -11 = b So, our finished equation is: y = 5x - 11!

OpenStudy (alexandrad):

thank you !